OpenStudy (anonymous):

The number of roots of the equation \(2 \sin^2\theta + 3\sin\theta + 1 = 0\) in the interval (0, 2) is

OpenStudy (anonymous):

its sin^2theta and sin theta

OpenStudy (anonymous):

(0,2pi)

OpenStudy (anonymous):

if we put 0 we get 1 if we put 2pi we get?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

x=sin theta then 2x^2+3x+1=0

OpenStudy (anonymous):

x=-0.5,-1

OpenStudy (anonymous):

sinx=-1/2,sinx=-1

ganeshie8 (ganeshie8):

yes!

ganeshie8 (ganeshie8):

solve each of them : sinx = -1/2 x = ?

ganeshie8 (ganeshie8):

what angles make the sin value equal to -1/2 ?

ganeshie8 (ganeshie8):

familiar with unit circle, right ?

OpenStudy (anonymous):

5pi/6

OpenStudy (anonymous):

unit circle means circle with radius 1

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

there will be two angles that make sin value equal to -1/2 :